Full Unanswered attempts leftSubm Assigned as Homework A well-thrown ball is caught in a well-padded ball first touches the mitt until it stops, what was the initiat velocity of the balr? (take the initial velocity of the ball to be positive) mitt. If the deceleration of the ball is 21429 m/s 2 and 2.4 ms elapses from the time the 4 attempts left 957 PM
Answer #1

Then given acceleration =a=21429 m/s^2

Time =t=2.4 ms

From the kinematics equation,

V=U+at

0=U-(21429)(2.4*10^-3)

0=U-51.43 m/s

U=51.43 m/s Answer

I hope help you !!